Events A and B are independent.
Find the indicated Probability.
P (A) = 0.39
P (B) = _______ (round answer to 3 decimal places).
P (A and B) = ) 0.02

Answer by math_tutor2020(3835)


Answer: 0.051

-------------------

Explanation:

We use the idea that
P(A and B) = P(A)*P(B)
which only applies if A and B are independent

Rearranging terms gets us
P(B) = P(A and B)/P(A)
P(B) = 0.02/0.39
P(B) = 0.05128205128206
P(B) = 0.051